OpenStudy (paki):

a person's remaining income after tax is 3000. tax rate is 25% .how much amount did he paid as tax.????

OpenStudy (hoblos):

let the total income be n when the tax takes tax take 25% of the income, 3000 are left thus n - 25% of n = 3000 can you find n now ?

OpenStudy (paki):

12000

OpenStudy (hoblos):

how did you get that ?

OpenStudy (paki):

i got that from your solution... am i rite...?

OpenStudy (paki):

n = 3000*100/25

OpenStudy (hoblos):

you are missing the first n in the equation it is n - (25/100)n = 3000

OpenStudy (paki):

then am i rite... n=3000+25/100

OpenStudy (hoblos):

you missed the second n now, you have to take n as a common factor n - (25/100)n = 3000 n(1 - 0.25) = 3000

OpenStudy (paki):

any good formula for solving such question...

OpenStudy (hoblos):

in other approach when we pay 25% of n, we are left with 3000 thus 3000 = 75% of n can you get n now ?

OpenStudy (paki):

yeah now got it... thanks :-)

OpenStudy (hoblos):

note that what you need is the tax, which is 25% of n

OpenStudy (paki):

hmm yeah, your second method is good...

OpenStudy (hoblos):

you can tell me what you get if you need to check your answer

OpenStudy (paki):

4000

OpenStudy (hoblos):

that's for n or for the tax ?

OpenStudy (paki):

foe n... for tax it will be 1000

OpenStudy (hoblos):

yeah, correct answer

OpenStudy (paki):

bundle of thanks for helping me :-)

OpenStudy (hoblos):

any time :)